Suppose P(A)=0.55, P(B)=0.45, P(A∪B')=0.7.

Find P(A∩B). Enter your answer to 3 decimal places.
Answer #1

P(AUB) = P(A) + P(B) - P(ANB)

  = P(A) + P(B) - P(A) + P(ANB)

= P B) + PAN B)

=1-PB) + P(ANB)

Substitute here known values

0.7 = 1-0.45 + P( A B )

=> 0.7 = 0.55 + P(ANB)

=> P(ANB) = 0.150